# A Survey on Integrated Sensing, Communication, and Computation

### v1

@@ VII-C Space-Air-Ground Integrated Networks Supported ISCC
By integrating satellite, aerial, and ground communication systems, SAGIN has been a promising solution for providing infrastructure-free environments with communication and intelligent services such as precision agriculture, environmental monitoring, and wildlife monitoring [15]. On one hand, compared with conventional terrestrial networks, SAGIN can provide global coverage especially for remote areas by using UAVs as relays and links between UAVs and satellites for providing backhaul connectivity. On the other hand, satellites are equipped with many remote sensors for detecting and monitoring the characteristics of earth systems from the space in a global perspective. This enables real-time and accurate decision-making for many global mission critical such as disaster assessment. Based on the satellites’ two types of functions mentioned above, two SAGIN-supported ISCC frameworks are introduced below.
@@
VII-C1 Terrestrial Sensing based ISCC framework in SAGINs
Consider a hierarchical SAGIN as shown in Fig. 12, where UAVs relay the sensory data from many terrestrial sensing devices to satellites for completing real-time intelligent tasks. Since computation migration between different satellites is difficult due to the high communication load and weak channel link gains between different satellites, only one satellite is selected for completing one task. Each UAV accesses several sensing devices within a disjoint coverage. The communication links between UAVs and the terrestrial sensing devices in their coverage are strong but the links between UAVs and satellites are weak due to the large path loss caused by the long distance. To tackle this issue, all UAVs collect sensory data from the sensing devices in their own coverages, pre-process the collected data for feature extraction and compression to save communication overhead, and forward the extracted and compressed feature. The satellite conducts a global fusion of the extracted and compressed features from all UAVs for executing the task. Compared with traditional terrestrial networks, SAGINs have the ability to connect the sensing devices in remote areas and can support much larger sensing ranges for more complicated ISCC applications.
As well as dealing with the high coupling among sensing, communication, and computation, there are several novel challenges for implementing the above SAGIN-supported ISCC framework. To begin with, satellites have high mobility and the path loss between UAVs and satellites is large, leading to weak and varying communication channels. As a result, new communication techniques are desirable. Next, a satellite has high coverage, resulting in large distance variations between the satellite and different connected UAVs. Consequently, communication time synchronization turns to be a severe issue. Besides, the issues of UAV trajectory design and satellite selection are also significant. Hence, the exploration of such a framework remains an open issue and calls for further efforts.
@@ VII-C1 Terrestrial Sensing based ISCC framework in SAGINs
VII-C2 Remote Sensing based ISCC Framework in SAGINs
@@
Another main function of satellites is remote sensing by being equipped with different types of sensors. As shown in Fig. 13, by obtaining different modalities of sensory data such as panchromatic images, multispectral images, thermal imaging infrared data, and radar sensory data, and transmitting the processed features to the ground station, many real-time tasks can be accomplished such as precision agriculture and climate change detection. To establish such systems, joint design of multi-modal sensing across multiple satellites, on-satellite processing, and satellite-ground communications are required. As well as addressing the same challenge of weak and fast varying channel links as the terrestrial-sensing-based ISCC framework, there are new design requirements including satellite scheduling for selecting the best sensing views and areas of the task, the multi-modal data alignment across different satellites, and the multi-modal data fusion schemes. For example, the data fusion schemes can be categorized into three types. The first is that satellites directly transmit the raw sensory data to the ground station. This scheme can keep the most sensing information but encounters severe communication bottlenecks for transmitting high-resolution sensory data. The second scheme is called early fusion, where each satellite transmits an extracted low-dimensional feature map to the ground station by saving communication load but sacrificing a part of the information. The last one is late fusion, where on-satellite decisions are made and only the results are fused at the ground station via e.g., a majority vote. In this scheme, only very little information is shared to the ground station but high communication efficiency can be achieved. Besides, since late data fusion shares minimum information to the ground station, it can be conducted when satellites and the ground station belong to multiple institutions.
VIII Conclusion
@@
ISCC, with the ability to improve network resource utility and achieve customized goals of complicated tasks, turns to be a key technology for realizing connected intelligence in future networks. In this paper, we first provide a comprehensive survey in terms of the historical partial integration technologies, i.e., ICC, ISC, and ISAC. Then, a holistic discussion on the shortages of these technologies is conducted to motivate ISCC, including low resource utilization, mismatch between goals of the overall task and each module (i.e., sensing, communication, or computation), and ignoring the tight coupling of different modules. Subsequently, the new challenges of ISCC is discussed, including inference management, high signal design complexity, complicated resource management, and lack of unified design criteria. To tackle these challenges, the techniques of signal design and network resource management for ISCC are surveyed. Finally, the implementation of ISCC in several future advanced networks is investigated. Hopefully, this article will promote the advancement of ISCC in 6G, encouraging future research efforts in this interesting field.

### v2

@@ VIII-C Space-Air-Ground Integrated Networks Supported ISCC
By integrating satellite, aerial, and ground communication systems, SAGIN has been a promising solution for providing infrastructure-free environments with communication and intelligent services such as precision agriculture, environmental monitoring, and wildlife monitoring [16, 332]. On one hand, compared with conventional terrestrial networks, SAGIN can provide global coverage especially for remote areas by using UAVs as relays and links between UAVs and satellites for providing backhaul connectivity. On the other hand, satellites are equipped with many remote sensors for detecting and monitoring the characteristics of earth systems from the space in a global perspective. This enables real-time and accurate decision-making for many global mission critical such as disaster assessment. Based on the satellites’ two types of functions mentioned above, two SAGIN-supported ISCC frameworks are introduced below.
@@
VIII-C1 Terrestrial Sensing based ISCC framework in SAGINs
Consider a hierarchical SAGIN as shown in Fig. 14, where UAVs relay the sensory data from many terrestrial sensing devices to satellites for completing real-time intelligent tasks. Since computation migration between different satellites is difficult due to the high communication load and weak channel link gains between different satellites, only one satellite is selected for completing one task. Each UAV accesses several sensing devices within a disjoint coverage. The communication links between UAVs and the terrestrial sensing devices in their coverage are strong but the links between UAVs and satellites are weak due to the large path loss caused by the long distance. To tackle this issue, all UAVs collect sensory data from the sensing devices in their own coverages, pre-process the collected data for feature extraction and compression to save communication overhead, and forward the extracted and compressed feature. The satellite conducts a global fusion of the extracted and compressed features from all UAVs for executing the task. Compared with traditional terrestrial networks, SAGINs have the ability to connect the sensing devices in remote areas and can support much larger sensing ranges for more complicated ISCC applications. In this framework, ISCC designs are demanded to jointly coordinate the terrestrial sensing, communication from sensors to UAVs, signal pre-processing on UAVs, communication between UAVs to the satellite, and the on-satellite computation to achieve high task performance.
@@ VIII-C1 Terrestrial Sensing based ISCC framework in SAGINs
VIII-C2 Remote Sensing based ISCC Framework in SAGINs
@@
Another main function of satellites is remote sensing by being equipped with different types of sensors. As shown in Fig. 15, by obtaining different modalities of sensory data such as panchromatic images, multispectral images, thermal imaging infrared data, and radar sensory data, and transmitting the processed features to the ground station, many real-time tasks can be accomplished such as precision agriculture and climate change detection. To establish such systems, joint design of multi-modal sensing across multiple satellites, on-satellite computation, and satellite-ground communications are required. As well as addressing the same challenge of weak and fast varying channel links as the terrestrial-sensing-based ISCC framework, there are new design requirements including satellite scheduling for selecting the best sensing views and areas of the task, the multi-modal data alignment across different satellites, and the multi-modal data fusion schemes. For example, the data fusion schemes can be categorized into three types. The first is that satellites directly transmit the raw sensory data to the ground station. This scheme can keep the most sensing information but encounters severe communication bottlenecks for transmitting high-resolution sensory data. The second scheme is called early fusion, where each satellite transmits an extracted low-dimensional feature map to the ground station by saving communication load but sacrificing a part of the information. The last one is late fusion, where on-satellite decisions are made and only the results are fused at the ground station via e.g., a majority vote. In this scheme, only very little information is shared to the ground station but high communication efficiency can be achieved. Besides, since late data fusion shares minimum information to the ground station, it can be conducted when satellites and the ground station belong to multiple institutions.
VIII-C3 Challenges
The implementation of SAGIN-supported ISCC faces several paractical challenges as well as that mentioned in Section V-C. First, the long distance and environmental issues, such as bad weather conditions, atmospheric disturbances, electromagnetic interference, and physical obstacles may degrade the communication and remote sensing performance. Second, satellites have high mobility and the path loss between UAVs and satellites is large, leading to weak and varying communication channels. Next, each satellite has a wide coverage, resulting in large distance variations between the satellite and different connected UAVs. Consequently, communication time synchronization becomes a severe issue. As well, the battery life of UAVs and ground devices may limit their capabilities of sensing, communication, and computation. Moreover, the SAGIN has a large coverage to support a massive number of devices. Coordinating the sensing, communication, and computation processes for them is critically challenging. To this end, it calls for future research efforts to implement SAGIN-supported ISCC.
VIII-D Discussion of Unresolved Issues
Although many efforts have been made toward ISCC as elaborated in Sections VI and VII, there are many critical issues remaining unresolved to implement ISCC in the future. Some of them are discussed below.
  • Lack of Theoretical Analysis Tools: Although there are individual theoretical analysis for sensing, communication, and computation, and their partial integration e.g., ISAC and AirComp, there is a lack of theory that can characterize the integration performance of the three modules, especially from the information-theoretic perspective. This further results in a lack of guidance to design common ISCC strategies that can be generally applied in different tasks, scenarios, and networks.
  • Hardware Design: New hardware designs are required to enable resource sharing and coordination among sensing, communication, and computation modules. As an example, in an ISCC-supported online federated learning task, a pipeline design, where on-device computation can be conducted to execute stochastic gradient descent right after obtaining each sensory data sample, can enhance the time efficiency but calls for dedicated hardware that breaks up the barriers between sensing and computation.
  • Protocols and Standardization: Although many schemes and algorithms are proposed for ISCC, the designs of ISCC protocols and standardization remain uncharted. Particularly, an important issue is that the ISCC protocol should be the compatible with existing wireless communication systems.
  • Security: In some critical tasks, it is enssential to avoid information leakage during sensing, communication, and computation processes. However, the ISCC design enhances the correlation between the three processes. For example, it is possible to infer the sensory data information based on the communication and computation loads and the computational results. It is also possible to infer the computational task information according to the sensory view and modalities. As a result, more advanced security methods are needed in ISCC systems.
  • Privacy: In ISCC systems, the completion of a task generally requires sharing sensory of intermediate data between different devices, edge servers, or even clouds. Data privacy preservation become an important issue. On one hand, we can utilize the distortion in the ISCC systems, e.g., sensing/communication noise and quantization error, to preserve data privacy. On the other hand, it is challenging to preserve data privacy while guaranteeing good task performance.
  • Robustness: ISCC systems should be robust to unexpected device failure and the case where some functions (e.g., sensing or communication) of one or several devices get lost. Novel techniques are required to guarantee the robustness of ISCC systems. Besides, similar to the privacy issue, robustness enhancement may need the sacrifice of task performance. Balancing the tradeoff among the robustness, privacy, task performance, and other issues is important in the practical implementation of ISCC.
  • Backward Compatibility with Existing Technologies: ISCC-enabled wireless systems should be compatible with existing wireless technologies. They should support old protocols and devices equipped with old hardware.
Overall, the efficient implementation of ISCC calls for many future research efforts to address the unresolved issues listed above.
IX Conclusion
@@
ISCC, with the ability to improve network resource utility and achieve customized goals of complicated tasks, turns to be a key technology for realizing connected intelligence in future networks. In this paper, we first provide a comprehensive survey in terms of the historical partial integration technologies, i.e., ICC, ISC, and ISAC. A holistic discussion on the shortages of these technologies is conducted to motivate ISCC, including low resource utilization, mismatch between goals of the overall task and each module (i.e., sensing, communication, or computation), and ignoring the tight coupling of different modules. Then, ISCC’s benefits, functions in 6G for supporting the KPIs and usage scenarios, and challenges of ISCC are analyzed. Particularly, the benefits include sensing-communication-computation symbiosis, the ability to achieve higher network resource utilization and task performance. The challenges include multi-objective optimization, inference management, high signal design complexity, complicated resource management, lack of unified design criteria, and generalization. To tackle these challenges, the techniques of signal design and network resource management for ISCC are surveyed. Finally, the implementation of ISCC in several future advanced networks is investigated and the unresolved issues which call for future research efforts are discussed. Hopefully, this article will promote the advancement of ISCC in 6G, encouraging future research efforts in this interesting field.